write 7x^2+5=x^2-8x in STANDARD form and find the value of b.
Answer by stanbon(75887)

write 7x^2+5=x^2-8x in STANDARD form and find the value of b.
----------------
6x^2 + 8x + 5 = 0
b = 8
========================
Cheers,
Stan H.